How do you express log4+3loga-1/2logb as a single logarithm?
Answer by user_dude2008(1861)

log4+3loga-1/2logb = log(4)+log(a^3)-log(b^(1/2)) = log(4)+log(a^3)-log(sqrt(b)) = log[(4a^3)/sqrt(b)]